OpenStudy (shesolitt):

Simplify open parentheses x to the 2 ninths power close parentheses to the 3 eighths power. x to the 5 seventeenths power x to the 1 twelfth power x to the 11 over 72 power x to the 43 over 72 power

OpenStudy (shesolitt):

I WILL MEDAL AND FAN

OpenStudy (shesolitt):

@mathmate

OpenStudy (mathmate):

Law of exponents: \(\large(x^a)^b=x^{ab}\) In your case, x=x a=2/9 b=3/8 Can you multiply fractions? If you need a review, here it is: https://www.mathsisfun.com/fractions_multiplication.html

OpenStudy (shesolitt):

im a little confused. 3/8 isnt a option @mathmate

OpenStudy (mathmate):

@shesolitt 3/8 should not be an option, nor 2/9! Please reread the second line of my previous post, carefully.

OpenStudy (shesolitt):

i got 1 over 12 @mathmate

OpenStudy (mathmate):

Good job, the first step is done! But...1/12 is not going to be the answer, although this is related to the answer. Reread the second line above, and see if you can find what's missing. Second line reproduced below: \(\large(x^a)^b=x^{ab}\)